### v1

# Robust and Resilient Networks with Integrated Sensing, Communication and Computation

Abstract
Emerging applications such as networked robotics, intelligent transportation, smart factories, and virtual and augmented reality demand integrated perception and connectivity enabled by wireless communication. This has driven growing interests in integrated sensing, communication, and computation (ISCC) systems, with a primary focus on their efficient co-designs. However, as ISCC systems increasingly support critical applications, they must not only deliver high performance but also demonstrate robustness and resilience. In this context, robustness refers to a system’s ability to maintain performance under uncertainties, while resilience denotes its capacity to sustain a minimum level of service in the face of major disruptions. To address this gap, this article presents an overview of ISCC systems from the perspectives of robustness and resilience under limited resources. First, key concepts related to these properties are introduced in the ISCC context. Subsequently, design approaches for realizing robust and resilient ISCC networks are discussed. Finally, the article concludes with the discussions of a case study and open research problems in this area.
I Introduction
Sensing and communications through wireless devices and networks have evolved largely independently over the past decades, while each playing a vital role in enabling new wireless services. Recently, the emergence of advanced applications in the sixth-generation (6G) wireless networks, such as intelligent healthcare, Industry 5.0, autonomous transportation, and smart cities, has created a pressing need for these two functions to work jointly. This has motivated intense studies on integrated sensing and communication (ISAC) in past years [1]. The relevant standardization progress for ISAC has been launched by 3GPP within the scope of IMT-2030 [2, 3].
Yet, the demands of these next-generation applications extend beyond the joint operation of sensing and communication. In the age of artificial intelligence (AI), networks must not only perceive and connect, but also reason and act intelligently. This then calls for edge-AI, which brings the computational capability closer to where data is generated [4]. However, the effectiveness of AI-based decision-making highly rely on the timely acquisition of environmental information and their delivery to the AI models. Therefore, the true intelligence comes from the integration of sensing, computing, and communication (ISCC), which allows applications to efficiently utilize the sensing information along with communication and computation resources to make smart decisions and reactions within a short time [5].
While 6G delivers more intelligent and pervasive ISCC services to the society, the underlying network architecture is becoming increasingly complex to support this interconnected set of functionalities. This complexity can at the same time lead to a more sensitive and vulnerable network, where disruptions and failures in any of the sensing, communication, and computing components could lead to service interruptions and, consequently, severe societal impacts [5]. In this context, the robustness and resilience aspects of the network have started to draw attention [6, 7]. “Robustness” refers to the ability of the system to tolerate a certain level of uncertainty without experiencing substantial performance degradation, that is the system should maintain its functionality within a pre-defined range of uncertainty. In contrast, “resilience” refers to the capability of systems to maintain a minimum performance even facing major disturbances, such as technical failures, human errors, jamming attacks or other unexpected strikes/changes, as well as to recover from these disruptions to an acceptable operational state over time [6].
While robustness and resilience of wireless networks have started to receive attentions, they remain underexplored in the context of the ISCC network [5]. In fact, these aspects are particularly critical for ISCC networks, which require careful orchestration of sensing, computation, and communication components to fulfill complex tasks. Such coordination not only include the common power, bandwidth, and computational resource management, but also asks for the tight synergy between different types of wireless nodes across different tiers of ISCC networks. Furthermore, effective operation relies on seamless coordination among software and hardware modules. These interdependencies create a highly complex network that is inherently sensitive and vulnerable to uncertainties and disruptions.
Figure 1: Exemplary key enablers of R2-ISCC networks.
To address these issues, this article advocates the development of robust and resilient ISCC (R2-ISCC) networks and discusses key design principles to achieve them. We first introduce the fundamental frameworks for ISCC resource management and provide the related critical concepts and metrics for characterizing robustness and resilience in these networks. Then, promising approaches from resource management, coding, architecture design, digital twins (DTs), and network perspectives that can provide robustness and resilience to ISCC networks are identified and discussed. In particular, we elaborate the fundamental ideas underpinning these capabilities and discuss how they interrelate with various enabling technologies. A case study on distributed sensing systems is presented to exemplify the design of R2-ISCC networks with respect to both robustness and resilience. Finally, we highlight key research directions and open problems for advancing R2-ISCC 6G networks.
II Overview of R2-ISCC Systems
This section first presents a common framework for resource management in ISCC systems and their response to uncertainties and disruptions. We then discuss how robustness and resilience are measured in R2-ISCC systems.
II-A Common Framework for R2-ISCC
Joint design in R2-ISCC: Joint design and resource management (RM) in ISCC networks aims to coordinate sensing, communication, and computing resources across devices, edge nodes, and the cloud to support diverse application tasks. This unified approach improves overall efficiency by treating resource provisioning as an integrated problem rather than managing each domain separately. Moreover, the joint design can be specialized to meet specific task requirements [8] via adaptive, goal-driven resource allocation.
While the integration of sensing, communication, and computing already increases the complexity of network operations and RM compared to traditional networks, the additional demand for precise and timely responses in critical applications further amplifies the ISCC network’s vulnerability to uncertainties and disruptions. These challenges may arise from the environmental randomness of wireless channels, (e.g., fading, interference, and noise), sensing and channel estimation errors, network dynamics, and intentional delays in distributed computing (e.g., stragglers). In addition, the internal disruptions (e.g., unexpected failures and outages) and external attacks (e.g., intentional and unintentional jamming and physical strikes) can further compromise network functionality and service availability. Therefore, designing R2-ISCC networks, which are robust and resilient to such uncertainties and disruptive events is urgently in-need. Fig. 1 illustrates several key enablers of R2-ISCC architectures, which will be discussed in detail throughout this article.
A continuous spectrum from robustness to resilience: Robust schemes are designed to handle uncertainties, typically using statistical models or worst-case bounds. However, they are inefficient for rare and high-impact events that require a shift in operational mode. Resilient schemes, by contrast, focus on responding to such disruptions when they occur. As illustrated in Fig. 2, the transition from robust to resilient design is not sharp as event rarity and impact lie on a continuum and are partly shaped by the designer’s perspective.
To accurately characterize the robustness and resilience of the network based on its state, the network phases illustrated in Fig. 3 are considered. Specifically, the time interval between $t_{0}$ and $t_{1}$ represents a normal operating phase, during which routine uncertainties (such as channel fading, expected network errors, system imperfections, and computational stragglers) may occur. These uncertainties can affect system performance, though not severely enough to disrupt operations. In this phase, the ISCC network must account for the impact of such uncertainties and minimize possible performance degradation by the means of robust designs and optimization techniques.
When the network begins to experience unexpected failures/disruptions that can drastically change its state, it enters the phase between $t_{1}$ and $t_{4}$. Basically, the onset of $t_{1}$, the network performance could quickly degrade due to these disruptions, whereby without any recovery mechanisms or intervention, the network may continue to operate at a severely diminished performance level. However, if resilient strategies are in place, the network can begin to recover either fully to its normal state or at least improve from the degraded condition as illustrated during the period from $t_{3}$ and $t_{4}$. It is important to note that resilience in network design encompasses more than just recovery. It also includes mechanisms that resist disruptions proactively, limit the extent of performance degradation during disruptions, and accelerate the recovery process [6]. These capabilities collectively contribute to the overall resilience of the ISCC network.
Figure 2: Illustration of the continuous spectrum from robust to resilient designs, based on the rareness of disruption events and the associated system states.
II-B Evaluation of Robustness and Resilience
To evaluate the robustness and resilience of an ISCC network, several quantitative metrics are commonly used. For robustness, key indicators include outage probability and service reliability under uncertain parameters. For instance, when latency is used as a QoS metric, the outage probability can be defined as the likelihood that a task fails to meet its latency requirement given that certain conditions (such as the mean user traffic rate) are only known with limited accuracy. Other performance metrics can also be applied to assess specific aspects of the ISCC network. For example, packet and bit error rates are useful for evaluating communication performance, while detection and false alarm rates serve as indicators of sensing performance.
For resilience, additional compound metrics are required to capture the inherent complexity of ISCC networks and their transitions through multiple operational phases during a disruption, as illustrated in Fig. 3.
In Phase 1, where performance begins to degrade, resilience can be evaluated based on the system’s resistance capability – its ability to maintain performance despite the onset of a disruption. One useful metric in this phase is the ratio between the time-averaged network performance under failure and that under normal conditions. Another useful metric is the worst-case performance during the disruption and service level fluctuations, which reflect how severely and variably the system degrades.
In Phase 2, where the network operates in a degraded state, the duration between $t_{2}$ and $t_{3}$ becomes important. This interval reflects the failure detection and response initiation time, a key indicator of responsiveness. Combined with Phase 3 (from $t_{3}$ and $t_{4}$), the total time from failure detection to system recovery ($t_{2}$ to $t_{4}$) defines the time-to-recover, a critical metric for measuring the overall resilience. Performance degradation during this interval can be further assessed using the delay in restoring system performance.
In some cases, full recovery to pre-failure performance levels may not be possible. Therefore, the post-recovery performance gap (the difference between the recovered and original performance levels) is also an important measure. Finally, over longer time spans, the mean time to failure (MTTF) should be considered to evaluate the long-term resilience of the network.
Figure 3: Illustration of different phases in an ISCC network with regard to robustness and resilience. This figure underlines that robustness and resilience are intertwined and characterize the system’s capability to withstand a spectrum of potential uncertainties and disruptions.
III Approaches for Realizing R2-ISCC Systems
This section outlines key principles for realizing R2-ISCC systems, followed by concrete enabling strategies.
III-A General Principles and Concepts
In the following, we introduce three classes of key principles that enable R2-ISCC, as shown in Fig. 4.
Passive strategies: From an RM perspective, diversity and redundancy are two fundamental concepts. Diversity ensures that multiple independent resources for communication, sensing, or computation are available to perform essential functions, which allows for fallback options when a component fails. Redundancy involves provisioning more resources than necessary, and enables the network to absorb unexpected events and potentially recover from failures. Robustness and resilience also require a shift in optimization focus beyond average-case performance to account for rare but impactful disruptions. This includes shaping the tail distribution of performance metrics to ensure the network can withstand the worst-case scenarios.
Proactive strategies: While passive strategies are applied by default for responding passively to disruptions, proactive strategies aim to enhance robustness and resilience by anticipating disruptions before they occur. When uncertainties, threats, or failures can be accurately predicted, the network can take preemptive actions to mitigate or avoid performance degradation. In fact, these strategies are increasingly feasible with integrated sensing capabilities. For example, if object tracking detects a potential blockage of a communication link (e.g., a building obstructing a beam), the transmitter can proactively adjust its beamforming direction to maintain connectivity and avoid service interruption.
Active strategies: In ISCC networks, when failures occur across sensing, communication, or computing components, protection and recovery mechanisms are essential for maintaining service continuity. Resilience depends on adaptive reconfiguration such as reassigning sensing tasks, rerouting communication links, or relocating computation to alternate nodes. Given limited resources, critical tasks must be prioritized, which demands for a highly reconfigurable ISCC architecture. Furthermore, learning from past disruptions enables the ISCC system to refine future preemption, mitigation, and recovery strategies, through a feedback loop that enhances long-term robustness and resilience.
Based on the aforementioned general concepts, the following discussion explores key techniques for realizing them within the context of ISCC systems.
Figure 4: Three classes of concepts for realizing R2-ISCC systems. Concrete examples of these strategies are presented in Subsections III-B to III-F.
III-B Robust and Proactive Resource Management
Robust RM: To enhance robustness in ISCC networks, a widely adopted approach is robust resource optimization, which explicitly accounts for various network uncertainties, including target variability, channel fluctuations, computing workloads, and resource availability. The core principle is to introduce redundancy and diversity, provisioning excess resources to safeguard against worst-case scenarios. For instance, multiple base stations (BSs), edge servers, or sensors may be assigned to the same task, Thusm these tasks can be dynamically redirected to alternative resource providers with more favorable conditions. While this ensures task completion under adverse conditions, it often results in inefficient resource use, as excessive provisioning is rarely needed in typical cases. To address this, advanced robust optimization methods have been developed that focus on the trade-off between average performance and outage probability. These approaches aim to balance high efficiency and strong reliability under uncertainty.
Proactive RM: Even when offering a good efficiency–reliability trade-off, robust approaches tend to provision an excessively large amount of resources because they lack precise knowledge of future conditions. In other words, to ensure reliability, they must account for a wide range of possible scenarios, many of which may never occur. However, if uncertain situations can be more accurately predicted, proactive and targeted resource management becomes possible, allowing the network to focus on a narrower uncertainty region and significantly improve resource utilization. Furthermore, traditional robust designs operate within a predefined tolerance level for fluctuations in network state. When disruptions exceed this threshold, the network can easily fail. To enhance resilience, the network must not only respond to failures but also anticipate severe events. With predictive capabilities, the network can proactively allocate resources in preparation. For instance, it could preemptively activate backup servers or reserve additional capacity before a disruption occurs.
III-C Coded Transmissions and Computations
By incorporating redundancy and diversity, the robustness and resilience of ISCC networks can be significantly enhanced. However, efficiently implementing these features is crucial, ideally achieving an optimal trade-off between average performance and reliability. To this end, the adoption of coded transmissions and coded computing in the network naturally arises. Channel coding is well-known for mitigating adverse communication conditions, while coded computing addresses the straggler problem often encountered in distributed computations [9]. Beyond applying coding techniques independently to communication and computation, the integrated nature of ISCC systems calls for joint coding approaches that account for dependencies across sensing, communication, and computation. Such holistic designs can further enhance overall performance and reliability despite potential disturbances affecting individual sub-systems. From a network-wide perspective, resilience requires that the system continues to function even when certain sensing, communication, or computation nodes fail. In this context, network-level coding enables data and function recovery, allowing the network to self-repair and maintain operations despite node outages.
III-D Distributed Multi-Tier Multi-Modal Network Architectures
Distributed sensing: In contrast to exploiting the judiciously designed resource management and coding schemes to provide robustness and resilience, another natural approach is to address these goals from the perspective of network architecture. Indeed, distributed network architecture inherently offers more robustness and resilience. This is because when sensing, communication, and computation functions are distributed across multiple nodes, the failure or degradation of any single node can be more easily compensated by others. Moreover, the decentralized nature of such architectures ensures that disruptions remain localized and less likely to cascade across the network.
Multi-modal sensing: In addition to spatial distribution, the use of multi-modal sensing, where multiple types of sensing modalities (e.g., camera, radar, lidar) are employed in parallel, can further enhance robustness and resilience of ISCC networks [10]. This is because performance degradation or functional failures often affect only specific types of sensors under certain conditions. For example, sun glare may impair camera-based sensing, while radar remains unaffected by lighting conditions. Conversely, radar may struggle with precise vertical motion detection, but cameras can handle more effectively. By incorporating complementary sensing modalities, the system can reduce the likelihood of falling into worst-case scenarios caused by single-modality failures.
Multi-tier sensing network: Extending the above concepts, the adoption of multi-tier and multi-layer network architectures introduces architectural redundancy and diversity, that further strengthen network robustness and resilience. For example, in dense small-cell deployments, a user can connect to multiple BSs to ensure continued service even if one BS fails. Moreover, with the evolution of wireless networks, services can now be supported by heterogeneous sources with diverse channel and system characteristics including ground-based BSs, unmanned aerial vehicles (UAVs), and low earth orbit (LEO) satellites. Each of these platforms offers unique strengths in terms of connectivity, sensing, and computing. By complementing one another, they collectively enhance the network’s ability to maintain functionality under varying conditions and disruptions [11].
III-E Digital Twin Assisted ISCC Networks
DT for R2-ISCC: The basic idea behind DTs is to create a virtual replica of the physical environment, that enable real-time simulation and analysis of system behaviors. In the context of ISCC networks, DTs can play a crucial role in predicting network states and anticipating real-world events [12], which makes them naturally suitable for enhancing preemption-based robustness and resilience. For example, if a DT predicts power outage at a BS, the network can preemptively offload traffic from the affected BS to neighboring ones, effectively minimizing service disruption. Beyond prediction, DTs are valuable for informed decision-making during and after disruptive events. Continuing with the above example, a DT can incorporate information about operational BSs, historical traffic patterns, and current service demand. Exploiting this knowledge, the network can develop more effective protection and recovery strategies, and optimize its response actions and resource allocation. Therefore, DTs can enhance both the anticipation and reaction capabilities of R2-ISCC networks.
Robust and resilient DT: Besides DT enabling R2-ISCC, the DT itself must be robust and resilient. In fact, DTs represent a key application of ISCC networks, as they often rely on the collection of diverse sensor data via communication networks, which is then processed to construct and update the digital representation. This dependency leads to critical design decisions regarding where the DT is implemented and how it interacts with the ISCC network. For instance, relying solely on a cloud-based DT may be inadequate, as such an approach could introduce latency and create a single point of failure. To address these concerns, it is desirable to deploy DTs in a distributed fashion at the edge, closer to users and network elements. However, deploying DTs at the edge presents another challenge, namely supporting the computationally intensive operations that some DTs require. Edge nodes may lack the processing power needed for full-scale simulation, prediction, and real-time synchronization with physical systems. Thus, efficient, lightweight DT models, possibly supported by edge-cloud collaboration, must be developed to ensure practical and scalable deployment.
III-F Resilience-by-Design ISCC
A network is referred to as resilience-by-design (RBD) when, already during the design phase, resilience bottlenecks (i.e., the most vulnerable components of the network) are identified, and targeted resilience provisions (such as redundancy, adaptive coding, or reconfigurable modules) are systematically embedded into the architecture [6]. In ISCC networks, the sensing, communication, and computing components contribute jointly to service delivery (e.g., sensing accuracy), but their susceptibility to failure often varies. Given the cost and complexity of resilience mechanisms, a key challenge in RBD R2-ISCC is to identify the most critical bottleneck and prioritize both offline resource provisioning and online resource management accordingly. To mitigate cascading failures, isolation of sensing, communication, and computing resources is essential. Furthermore, an RBD R2-ISCC network must support dynamic reconfiguration and the ability to shift operational modes based on the availability of functioning resources. For instance, sensors may transmit high-quality raw data when high-capacity links (e.g., cellular) to an edge server is available. However, if that link fails but a low-capacity fallback link (e.g., LoRa) is provisioned, the affected sensors may need to adapt by reducing data quality, lowering transmission frequency, or even sending only selected locally processed results. Similarly, the inference algorithm at the edge server must be able to adapt to significant variations in received data or even to missing sensor inputs. An RBD R2-ISCC foresees such contingency operational modes during the design phase, and develops reconfiguration strategies to enable seamless transitions among them.
IV Case Study and Open Problems
In this section, an illustrative case for an R2-ISCC network is first presented, which is then followed by various open problems.
IV-A Case Study
As discussed in Section III, realizing an R2-ISCC requires a design that is both robust to parameter uncertainties and resilient to disruptions. To illustrate the interplay between robustness and resilience, we consider sensor selection and power adaptation in a distributed radar network [13]. Thereby, the optimal radar is chosen based on the target’s estimated location. Moreover, to meet a required sensing SNR $\gamma_{\mathrm{req}}$, sensor $i$ transmits with power $P_{i}(n)\propto\hat{d}_{i}^{4}(n)$ under a LoS mono-static radar model, where $\hat{d}_{i}(n)$ is the estimated distance to the target at time $n$. We assume a location error of up to $\epsilon_{\max}=0.5$ m and potential SNR drops due to blockages. Fig. 6 show the sensing SNR (normalized to $\gamma_{\mathrm{req}}$) over 200 epochs along the trajectory in Fig. 5, comparing four sensor-selection and power-adaptation strategies.
Baseline
The baseline scheme selects the sensor with the smallest estimated distance and sets transmit power based on this point estimate, ignoring location uncertainty and potential LoS blockages. As observed from Fig. 6, tracking fails when the target enters a shadowed zone. Even in free space, the SNR requirement is not consistently met due to underestimating the power needs under location uncertainty.
Robust
The robust design considers the worst-case target location for both sensor selection and power adaptation. This leads to consistently meeting the sensing requirement, (i.e., above 00 dB) along every unobstructed segment, at the expense of a higher average transmit power.
Resilient
The resilient design proactively responds to disruptions, such as blockages, by selecting sensors with an unobstructed view. Unlike the previous schemes, it largely avoids sharp SNR drops. Remaining drops occur when the target is actually in a shadowed region, but its estimated location lies outside, revealing a lack of robustness.
Robust & Resilient
This robust and resilient scheme accounts for both location uncertainty and blockages, respectively. As a result, the sensing SNR reliably remains above the target threshold throughout the entire trajectory.
The above example illustrates an R2-ISCC design under location uncertainty and blockage in a simplified setting. Real-world ISCC systems are far more complex, involving various uncertain parameters as well as both known and unknown disruptions. This highlights the need for further research on R2-ISCC, with key directions reviewed next.
Figure 5: Illustration of a distributed sensor setup for cooperative target tracking. The true target location is uncertain (within blue circles); the estimated location (flag) includes an uncertainty region (green circles). Moreover, sensor views may be obstructed illustrated by walls.
Figure 6: Time evolution of the mean SNR (dB relative to target SNR) for sensors S1 to S4 under standard, robust, resilient, and robust & resilient schemes. The window length of the rolling mean is 10 epochs. The colored circles indicate the instantaneous SNR and which sensor has been selected.
IV-B Open Problems
To complete this paper, some critical open problems for R2-ISCC are discussed below.
Efficient and scalable resource management for ISCC: Although widely used in conventional networks, diversity and redundancy schemes are typically designed for individual purposes, without jointly addressing sensing, computing, and communication. In the context of ISCC networks, new strategies that can efficiently integrate diversity and redundancy across all three components are needed, particularly for tasks that exhibit intrinsic dependencies among these functions. Moreover, the joint optimization of sensing, computing, and communication resources results in inherently high-dimensional and complex optimization problems. Hence, the development of efficient, scalable solution methods is critical to enable practical deployment and operation of robust and resilient ISCC systems.
Task-oriented and criticality-aware ISCC: ISCC networks must handle heterogeneous tasks with varying requirements and urgency levels simultaneously. To enhance robustness through proactive resource management, it is essential that the system is aware of these task-level differences when allocating resources. This awareness becomes even more critical during disruptions, when resource availability is constrained and careful prioritization is essential. One promising approach to address this issue is network slicing, which enables application-aware resource allocation. However, conventional network slicing primarily focuses on communication resources, which is insufficient in the ISCC context. Therefore, holistic slicing approaches that can simultaneously manage sensing, communication, and computation resources are required.
Joint source–channel–computing network coding schemes: To support network-level resilience, coding must extend beyond individual node types and retain the network functionality even when some sensing, communication, or computation nodes become unresponsive or fail. For example, coded computing techniques should be extended to the ISCC setting, accounting for e.g. sensor failures. Similarly, network coding, known for its resilience to link failures, eavesdropping, and jamming, offers a promising foundation for designing ISCC architectures capable of withstanding diverse and unpredictable network-level disruptions.
Synergies of multi-tier multi-modal ISCC networks: Distributed multi-tier, multi-modal ISCC networks promise greater robustness and resilience levels through complementary features for fault tolerance. However, research in both multi-modal and distributed ISCC networks is still in its early stages, requiring new architectures and design methodologies to integrate heterogeneous devices across different locations. Moreover, fusion techniques must support efficient representation alignment such that the feature spaces across different sensor types are aligned for effective cooperative perception. Furthermore, from a multi-task perspective, multiple perception and AI tasks should be designed to share a unified feature space. This would allow a single, jointly sensed data representation to support multiple tasks, significantly reducing communication overhead.
Resilience bottleneck in RBD ISCC networks: The RBD principle emphasizes embedding resilience capabilities during the design phase. In ISCC networks, resilience provisions across sensing, communication, and computing components are inherently heterogeneous and contribute differently to the end-to-end service resilience. As a result, efficient provisioning requires identifying resilience bottlenecks and strategically investing in the components that yield the greatest improvement in overall resilience. This challenge is further amplified in task-oriented designs, where optimal provisioning trade-offs depend heavily on specific task objectives. While the impact of resilience provisioning in individual subsystems has been studied, a holistic, joint approach to RBD in ISCC networks remains largely unexplored.
Self-organizing ISCC networks: A critical aspect of the RBD principle is reconfigurability based on available resilience measures. However, to hugely benefit from the network reconfigurability, corresponding strategies that support effective task redistribution through resource reallocation and network reconfiguration are required. Furthermore, while existing self-organizing networks (SONs) in communication systems have partially addressed this need, they do not account for the specific requirement of ISCC networks, which involve far more complex and tightly coupled sensing, computing, and communication operations [14].
V Conclusions
ISCC networks are essential for enabling intelligent perception in domains such as networked robotics, smart healthcare, Industry 5.0, and autonomous transportation. However, the interdependencies among sensing, computing, and communication components make them inherently complex and more vulnerable to uncertainties and disruptions. To address this, robustness and resilience must be embedded into their design. This article proposes an R2-ISCC framework, and outlining key principles and strategies across resource management, coding, architecture, DTs, and network-level design. A simulated case study highlights the importance and interplay of robustness and resilience. Key research problems are identified to guide future work. We hope this study inspires continued exploration and innovation toward R2-ISCC networks.